A curve is such that dy∕dx=√ (2x+5) and (2,5) is a point on the curve. What is the equation of the curve?
1
Expert's answer
2020-04-21T15:55:36-0400

Given

"\frac{d y }{d x}=\sqrt{( 2 x+5)}"

Separate variables

"\begin{aligned}
dy&= \sqrt{2x+5}dx\\
y&=\int \sqrt{2x+5}dx\\
&=\frac{1}{2}\int 2\sqrt{2x+5}dx\\
&=\frac{1}{2}\frac{2}{3} (2x+5)^{3/2}+c\\
&= \frac{1}{3}(2x+5)^{3/2}+c
\end{aligned}"

Since the curve passes through (2,5) , then

"\begin{aligned}
5&= \frac{1}{3}(4+5)^{3/2}+c\\
5&= 9+c\\
c&=-4
\end{aligned}"

Hence

"y= \frac{1}{3}(2x+5)^{3/2}-4"